An engine being tested for fuel consumption was operated at two fixed settings. The first setting results in a fuel consumption of 1 gallon every 12 minutes. The second setting results in a fuel consumption of 1 gallon every 15 minutes. The test ran 5 hours and 22 gallons of fuel were consumed. For how many minutes was the first test on the first setting?
Answer by stanbon(48558)

An engine being tested for fuel consumption was operated at two fixed settings. The first setting results in a fuel consumption of 1 gallon every 12 minutes. The second setting results in a fuel consumption of 1 gallon every 15 minutes. The test ran 5 hours and 22 gallons of fuel were consumed. For how many minutes was the first test on the first setting?
---------
Let "x" minutes be be running time for the 1st setting;
# of gallons consumed will be x/12
----------
Then "300-x" will be the running time for the 2nd setting:
# of gallons consumes will be (300-x)/15
----------
EQUATION:
gallons + gallons = 22
x/12 + (300-x)/15 = 22
Multiply thru by 60 to get:
5x + 1200-4x = 60*22
x=1320-1200
x=120 minutes (time the 1st setting was used)
==========
Cheers,
Stan H.